# PortraitTalk: Towards Customizable One-Shot Audio-to-Talking Face Generation

### v1

@@ Introduction
AnimateNet
Enforcing proper motion in talking face videos is crucial, as it enables the creation of realistic talking face videos. To this end, we split the animation enforcement process into two distinct phases: audio-to-motion and motion-to-frame generation. The audio-to-motion component is responsible for producing a sequence of facial movements (here, face landmarks) that accurately represent expected lip movements. We utilize an audio-to-motion generation module (Ye et al. 2023b), which employs a variational autoencoder alongside the HuBERT audio Transformer (Hsu et al. 2021). This module is equipped with a dilated convolutional encoder and decoder which enhances the robustness of feature extraction and the generation of extended motion sequences. Following this, the motion-to-frame generation phase is expected to produce realistic talking face videos that are aligned with facial landmarks derived from the audio-to-motion module. To facilitate this, we introduce AnimateNet, which is developed based on the principles of ControlNet (Zhang, Rao, and Agrawala 2023).
Accurate audio-lip synchronization, while necessary, is not sufficient for producing realistic talking face videos. Temporal consistency is vital for video generation, as it ensures smooth and coherent transitions among video frames. In addition, variability in head movements, attributable to the stochastic nature of diffusion models, can complicate the creation of consistent and smooth transitions in the talking person’s motion from frame to frame. Drawing inspiration from the previous studies (Tian et al. 2024; Stypułkowski, Vougioukas et al. 2024), we utilize the last generated frame from the previous step as the initial point for the subsequent frame generation and head placement guidance, thereby ensuring continuity and consistency across frames. Specifically, for each frame, we extract facial landmarks corresponding to the associated speech segment. Meanwhile, we retrieve the head placement of the previously generated frame and use it to guide the head placement of the current frame. For the initial frame, we retrieve and utilize the head placement from the reference image. These acquired facial landmarks and head placements are then processed through dedicated encoders to generate respective guidance embeddings for landmarks and head placements. These embeddings are subsequently integrated into AnimateNet via distinct cross-attention blocks. The processed animation information is injected into IdentityNet via cross-attention mechanisms. The final cross-attention output incorporates information related to identity, facial structure and temporal information across adjacent frames.
@@
AnimateNet utilizes a similar objective function as the original stable diffusion:
@@
where $c$ refers to the set of conditions, including identity features, facial landmarks, head placement guidance and temporal information. This architectural design enables PortraitTalk to concentrate on audio-lip synchronization, while preserving detailed facial characteristics across generated frames. It ensures temporal consistency and synchronizes head movements throughout the generated talking video.
Audio-Driven Facial Dynamics Score
Existing metrics in audio-to-talking face generation often focus on individual aspects of talking face video (e.g., visual quality (Hong and Xu 2023), audio-lip synchronization (Chung and Zisserman 2016)). There is a need for a comprehensive metric that evaluates realistic facial details, temporal alignment, and motion coherency. The Audio-Driven Facial Dynamics (ADFD) score addresses this by assessing both the spatial alignment and dynamic movement of facial landmarks over time, advancing beyond traditional metrics by incorporating these additional factors relative to the audio. The audio-driven facial dynamics score is formulated as:
@@
×$\quad\times w_{2}\left(\frac{\left(\mathbf{M}_{t}^{gen}\cdot\mathbf{M}_{t}^{gt}\right)+1}{2\|\mathbf{M}_{t}^{gen}\|\|\mathbf{M}_{t}^{gt}\|}\right)$ (3)
@@
where, $T$ represents the total number of frames in a video, $L_{i}^{gen}$ and $L_{i}^{gt}$ are arrays of generated and ground truth landmarks, respectively, at each time step $t$. $d$ is the maximum distance between two points in the frame, which is used for normalization. Motion vectors of the generated and ground truth landmarks at time step $t$ are represented by $M_{t}^{gen}$ and $M_{t}^{gt}$. To determine these motion vectors, we calculate the difference between two consecutive frames. $w_{1}$ and $w_{2}$ are the balancing parameters.
@@
The ADFD score has two main parts. The first quantifies the spatial alignment of facial landmarks between generated and ground-truth talking face videos by measuring the normalized Euclidean distance between the corresponding landmarks. The second part assesses the motion coherence and temporal consistency of the facial movements by calculating the cosine similarity of the motion vectors. This part evaluates the extent to which the direction and scale of facial movements in the generated video correspond with those in the ground truth over time. To accommodate specific research needs, ADFD can be fine-tuned using adjustable weights ($w_{1}$ and $w_{2}$), allowing one to prioritize certain aspects of the evaluation. Further, to ensure that all values reflect degrees of similarity on a non-negative scale, the range of the cosine similarity component has been adjusted from $[-1,1]$ to $[0,1]$. A high value in the ADFD score (close to $1$) signifies that the generated talking face video not only aligns well spatially but also demonstrates temporal consistency across frames.
Experimental Settings
Datasets and Evaluation Metrics
@@
We use the HDTF (Zhang et al. 2021) and MEAD(Wang et al. 2020) datasets in our study. The HDTF dataset contains high-quality videos of over 300 identities from YouTube. For training, we randomly select 6 videos, approximately $30,000$ frames in total. Subsequently, these videos are carefully reviewed to ensure the training set represents diverse gender and ethical variations. MEAD contains $60$ speakers with 8 different emotions. This dataset is captured in a laboratory setting and each emotion is captured under three different emotion intensity levels. In this paper, four randomly selected videos from MEAD are used for training. For the testing stage, similar to  Tan et al. (2024a), we randomly select $20\%$ of the HDTF dataset and choose speakers ‘M003’, ‘M009’, ‘M030’, and ‘W015’ from MEAD.
Following previous work (Shen et al. 2023; Hong and Xu 2023), we employ well-established metrics such as PSNR (Horé and Ziou 2010), SSIM (Wang et al. 2004), and FID (Heusel et al. 2017) to assess the face generation capability of our model relative to previous state-of-the-art methods. To further evaluate the accuracy of facial landmark placement and the smoothness of facial motion transitions between frames, we use LandMark Distance (LMD) (Chen et al. 2019b) and our novel Audio-Driven Facial Dynamics (ADFD) score. Additionally, we incorporate the SyncNet error metric to specifically measure the synchronization between audio and visual elements, ensuring that our model effectively aligns lip movements with spoken words, a critical factor for evaluating the generated talking face videos.
Implementation Details
Our framework, constructed on the basis of Stable Diffusion v1.5 (Rombach et al. 2021), is implemented using PyTorch. We train our model with the AdamW (Loshchilov and Hutter 2017) optimizer and set the batch size to $32$. The learning rate is set to $65e-$ 6. It should be noted that after generating facial landmarks from audio speech, we refine these landmarks to closely align with the outputs from MediaPipe (Ablavatski et al. 2020). This refinement significantly enhances the training of AnimateNet and positively influences its generalization capability. The training process encompasses a multi-stage approach. Initially, each component is trained independently. IdentityNet is finetuned for 10 epochs, followed by the training of AnimateNet (excluding temporal attention) for $100$ epochs. Subsequently, in the second stage, the entire framework undergoes integrated training for $50$ epochs. All training and testing phases are conducted on a single NVIDIA A100 GPU. The overall training takes approximately two weeks.
Experiments
Quantitative Results
Method MEAD (Wang et al. 2020) HDTF (Zhang et al. 2021)
@@
MakeItTalk (Zhou et al. 2020) 19.442 0.614 2.541/2.309 37.917 5.176 0.693 21.985 0.709 2.395/2.182 18.730 4.753 0.726
Wav2Lip (Prajwal et al. 2020) 19.875 0.633 1.438/2.138 44.510 8.774 0.584 22.323 0.727 1.759/2.002 22.397 9.032 0.699
Audio2Head (Wang et al. 2021) 18.764 0.586 2.053/2.293 27.236 6.494 0.725 21.608 0.702 1.983/2.060 29.385 7.076 0.835
SadTalker (Zhang et al. 2023) 19.042 0.606 2.038/2.335 39.308 7.065 0.761 21.701 0.702 1.995/2.147 14.261 7.414 0.726
IP-LAP (Zhong et al. 2023) 19.832 0.627 2.140/2.116 46.502 4.156 0.527 22.615 0.731 1.951/1.938 19.281 3.456 0.633
TalkLip (Wang et al. 2023) 19.492 0.623 1.951/2.204 41.066 5.724 0.548 22.241 0.730 1.976/1.937 23.850 1.076 0.518
EAMM (Ji et al. 2022) 18.867 0.610 2.543/2.413 31.268 1.762 0.847 19.866 0.626 2.910/2.937 41.200 4.445 0.839
EDTalk (Tan et al. 2024a) 21.628 0.722 1.537/1.290 17.698 8.115 0.715 25.156 0.811 1.676/1.315 13.785 7.642 0.722
@@
Table 1: A quantitative comparison with the SOTA methods on MEAD and HDTF. The highest scores for each metric are highlighted in bold. The symbols ‘↑’ and ‘↓’ denote that higher and lower values, respectively, indicate better performance.
@@
To assess our model, we conducted a quantitative comparison with the state-of-the-art methods in audio-to-talking-face generation. The results are reported in Table 1. According to the table, PortraitTalk consistently outperforms the existing approaches across nearly all the metrics on the HDTF and MEAD datasets. Wav2Lip (Prajwal et al. 2020) achieves a relatively higher SyncNet score on HDTF, primarily because it utilizes this metric as one of its training loss functions. Our model achieves a satisfactory SyncNet score, surpassing Wav2Lip on the MEAD dataset, and excels in audio-lip alignment and facial structure accuracy, as indicated by LMD and ADFD metrics. Although EAMM has the highest ADFD score, our model ranks closely behind on the HDTF dataset, showing similar effectiveness. The slightly larger gap on the MEAD dataset is likely due to its wider range of emotional expressions and varying intensities which directly impacts lip movements.
@@
Figure 3: A qualitative comparison with the existing talking face generation methods. Please note that the methods in the dash box use external emotion labels or reference videos to generate expressive videos.
Qualitative Results
@@
In Figure 3, we compare the generation quality of PortraitTalk against the state-of-the-art approaches. MakeItTalk and Wav2Lip produce low-quality outputs with artifacts, notably around the mouth, which fail to blend seamlessly with adjacent facial features. Audio2Head (Wang et al. 2021), TalkLip (Wang et al. 2023), and IP-LAP (Zhong et al. 2023) exhibit shortcomings in achieving precise audio-lip synchronization, with generated mouth movements often slightly open and displaying distortions across frames, thereby diminishing the realism of the output.
@@
In emotional talking face generation, existing methods typically generate videos based on discrete emotional labels or facial expressions extracted from user-provided reference-style videos. This approach, however, has significant limitations. EAMM (Ji et al. 2022) and EmoGen (Goyal et al. 2023), for example, struggle to maintain identity consistency across the generated videos. This issue is particularly visible in Figure 3, involving the male subject with EAMM and the female subject with EmoGen. Furthermore, EmoGen’s modifications are restricted to the square space of the facial region, leading to blurred areas in each frame. Such inconsistencies severely compromise the realism and visual quality of the generated talking face videos. EDTalk (Tan et al. 2024a) shows improved performance in generating emotional talking faces. However, it also exhibits noticeable jitter and irregular mouth shapes in the generated frames. In contrast, our model demonstrates superior performance in both emotion-agnostic and emotional talking face generation. It excels at preserving identity consistency, ensuring precise audio-lip synchronization, and maintaining consistent head movements throughout the video.
@@
Table 2: Quantitative ablation results.
Ablation Study
Figure 4: Visualization of the results obtained by including different components in PortraitTalk.
@@
Unified Attention Block: Initially, we merge the separate cross-attention blocks of IdentityNet into a single attention block to assess its impact on performance. The results are shown in Table 2 and Figure 4. The findings indicate a significant degradation in the model’s ability to preserve the intended identity, with its recognition capability reduced to identifying only the gender from text prompts.
@@
Impact of the Face Encoder: Further, we substitute our dedicated face encoder with a CLIP image encoder to explore the influence of the encoding mechanism on the generation capability of PortraitTalk. While there is an observable enhancement in image fidelity and identity consistency compared to the initial ablation scenario, the model still struggles accurately to maintain detailed identity attributes across frames. These experiments validate our decision to employ a specific face encoder optimized for facial feature extraction, underscoring its importance in achieving high-quality and consistent talking face generation.
@@
Number of Reference Images: We investigate the impact of using multiple reference images on the quality and realism of generated visual content. As demonstrated in Figure 4, employing multiple images of the same identity, showcasing various expressions or head angles, significantly enhances the model’s capability to learn the underlying facial structure of that identity. This strategy is essential for facilitating customization in generated talking face videos, enabling the model to effectively respond to a variety of customization prompts. In contrast, reliance on a single reference image leads to a considerable reduction in both, the quality and realism of the generated faces. This issue is especially evident in male case (Figure 4), where the model produces an overly oval head shape.
@@
Impact of Structure and Temporal Attention: We also try to remove the head placement guidance and temporal attention mechanisms from AnimateNet. This modification leads to a noticeable decrease in coherence across frames in the generated videos. As illustrated in Figure 4, while the lip synchronization remains somewhat accurate, there is a clear inconsistency in head placement and noticeable distortions between consecutive frames. Interestingly, without temporal attention, the generative model often fails to maintain consistency in the clothing of the generated talking faces. This variation underscores the crucial role of head placement guidance and temporal attention in maintaining the stability and continuity of head movements throughout the video. s
Figure 5: Diverse, expressive, and customizable talking faces generated by PortraitTalk.
Applications and Limitations

### v2

@@ III-A IdentityNet
III-B AnimateNet
Enforcing accurate facial motion is essential for generating realistic talking face videos. To address this, we divide the animation process into two phases: audio-to-motion and motion-to-frame generation. In the audio-to-motion phase, a sequence of facial movements (represented as facial landmarks) is generated to capture the expected lip and facial dynamics. This is achieved using an audio-to-motion module [27], which combines a variational autoencoder with the HuBERT audio Transformer [50]. The module employs a dilated convolutional encoder and decoder to improve the robustness of feature extraction and support the generation of longer, coherent motion sequences. Following generation, the facial motion representations are refined to align with MediaPipe [51] landmarks, further enhancing AnimateNet’s training stability and generalization across diverse speaking styles.
The motion-to-frame generation phase is responsible for producing realistic talking faces aligned with the motion features extracted from the audio. To achieve this, we introduce AnimateNet, built upon SD 1.5 and inspired by the design principles of ControlNet [52]. In ControlNet, an additional trainable branch is introduced alongside a frozen diffusion model to condition the generation on external inputs without disrupting the original generative capabilities. Similarly, AnimateNet integrates trainable control branches that modulate the generation process based on motion-related conditions, enabling the model to faithfully synthesize speech-driven facial movements. Motion features extracted by AnimateNet are fused with IdentityNet through ZeroConv layers, allowing seamless integration of motion and appearance information without disrupting the learned representations of either network. Additionally, AnimateNet employs three distinct cross-attention mechanisms (Structure, Identity, and Temporal), each designed to process a specific condition.
The structure cross-attention, processes facial landmarks corresponding to the associated audio segment, ensuring accurate alignment of audio-lip movements. Next, identity cross-attention focuses on head placement guidance, which is derived from the extracted head contour of the previous talking head or, for the initial frame, the reference image. To obtain this guidance, we extract the head contour from each frame using a modified Canny edge detector with reduced sensitivity to local details. This produces a coarse binary edge map that captures the overall shape and position of the head while ignoring fine-grained facial features. We encode the contour maps through a shallow convolutional branch and inject them into the diffusion via ZeroConv layers. This allows the network to condition the generation on head placement without interfering with identity or descriptive features. This approach addresses the challenge of variability in head movements, attributable to the stochastic nature of diffusion models, ensuring that head movements remain consistent and realistic across frames. To ensure smooth transitions between frames, and inspired by prior work [53, 37], we condition the generation at each time step on the frame produced in the previous step using a temporal cross-attention mechanism. Finally, a weighted sum of each attention block ($Z^{\text{new}}$) is computed and serves as the input for subsequent layers. The attention computation can be indicated as:
$Z^{\text{new}}=$ $\,w_{1}\left(\text{CrossAttn}_{\textbf{Structure}}(Q,K_{1},V_{1})\right)$
$+w_{2}\left(\text{CrossAttn}_{\textbf{Identity}}(Q,K_{2},V_{2})\right)$
$+w_{3}\left(\text{CrossAttn}_{\textbf{Temporal}}(Q,K_{3},V_{3})\right)$ (2)
where $\text{CrossAttn}(Q,K_{i},V_{i})$ computes the attention score following the standard attention [54] mechanism. $Q$, $K$ and $V$ represent the query, key and value matrices of the attention process. While the query remains shared across conditions, each condition-specific branch learns independent key-value mappings. The relative importance of each condition is controlled via learnable weights ($w_{i}$). The outputs of these attention mechanisms are integrated into IdentityNet, where the weighted sum of the resulting feature maps is used for the final generation step.
@@ III-B AnimateNet
For training, AnimateNet optimizes a denoising diffusion objective similar to stable diffusion:
@@
where $c$ refers to the set of conditions, including facial landmarks, head placement guidance and temporal information. By decoupling condition learning and leveraging specialized cross-attention mechanisms, AnimateNet achieves robust temporal consistency, natural head movements, and accurate audio-lip synchronization, resulting in realistic and identity-consistent talking face videos.
III-C Audio-Driven Facial Dynamics Score
Existing metrics in audio-to-talking face generation often focus on individual aspects of talking face video (e.g., visual quality [55], audio-lip synchronization [56]). There is a need for a comprehensive metric that evaluates realistic facial details, temporal alignment, and motion coherence. To this end, we advocate Audio-Driven Facial Dynamics (ADFD) that assesses both the spatial alignment and dynamic movement of facial landmarks over time, advancing beyond traditional metrics by incorporating these additional factors relative to the audio. The ADFD metric is formulated as:
@@ III-C Audio-Driven Facial Dynamics Score
$\quad\times w_{2}\left(\frac{\left(\mathbf{M}_{t}^{gen}\cdot\mathbf{M}_{t}^{gt}\right)+1}{2\|\mathbf{M}_{t}^{gen}\|\|\mathbf{M}_{t}^{gt}\|}\right)$ (4)
@@
where, $T$ represents the total number of frames in a video, $L_{i}^{gen}$ and $L_{i}^{gt}$ are the arrays of generated and ground truth landmarks, respectively, at each time step $t$. $d$ is the maximum distance between two points in the frame, which is used for normalization. Motion vectors of the generated and ground truth landmarks at time step $t$ are represented by $M_{t}^{gen}$ and $M_{t}^{gt}$. To determine these motion vectors, we calculate the difference between two consecutive frames. $w_{1}$ and $w_{2}$ are the balancing parameters.
@@
The ADFD score has two main parts. The first quantifies the spatial alignment of facial landmarks between generated and ground-truth talking face videos by measuring the normalized Euclidean distance between the corresponding landmarks. The second part assesses the motion coherence and temporal consistency of the facial movements by calculating the cosine similarity of the motion vectors. This part evaluates the extent to which the direction and scale of facial movements in the generated video correspond with those in the ground truth over time. To accommodate specific research needs, ADFD can be fine-tuned using adjustable weights ($w_{1}$ and $w_{2}$), allowing one to prioritize certain aspects of the evaluation. Further, to ensure that all values reflect degrees of similarity on a non-negative scale, the range of the cosine similarity component has been adjusted from [1,1][-1,1] to $[0,1]$. A high value in the ADFD score (close to $1$) signifies that the generated talking face video not only aligns well spatially but also demonstrates temporal consistency across frames.
IV Experimental Results
IV-A Datasets and Evaluation Metrics
@@
We use the HDTF [57] and MEAD[58] datasets for evaluation. The HDTF dataset contains high-quality videos of over 300 identities from YouTube. For training, we randomly select 6 videos, approximately $30,000$ frames in total. MEAD contains $60$ speakers with 8 different emotions. This dataset is captured in a laboratory setting and each emotion is captured under three different emotion intensity levels. In this paper, four randomly selected videos from MEAD are used for training. For the testing stage, similar to [2], we randomly select $20\%$ of the HDTF dataset and choose speakers ‘M003’, ‘M009’, ‘M030’, and ‘W015’ from MEAD. Since both HDTF and MEAD lack text prompts, we manually created descriptive prompts for the selected videos used in fine-tuning, capturing key attributes such as gender, background, and hair color.
Method MEAD [58] HDTF [57]
@@
MakeItTalk [5] 19.442 0.614 2.541/2.309 37.917 5.176 0.693 21.985 0.709 2.395/2.182 18.730 4.753 0.726
Wav2Lip [12] 19.875 0.633 1.438/2.138 44.510 8.774 0.584 22.323 0.727 1.759/2.002 22.397 9.032 0.699
Audio2Head [33] 18.764 0.586 2.053/2.293 27.236 6.494 0.725 21.608 0.702 1.983/2.060 29.385 7.076 0.835
SadTalker [34] 19.042 0.606 2.038/2.335 39.308 7.065 0.761 21.701 0.702 1.995/2.147 14.261 7.414 0.726
IP-LAP [35] 19.832 0.627 2.140/2.116 46.502 4.156 0.527 22.615 0.731 1.951/1.938 19.281 3.456 0.633
TalkLip [59] 19.492 0.623 1.951/2.204 41.066 5.724 0.548 22.241 0.730 1.976/1.937 23.850 1.076 0.518
EAMM [13] 18.867 0.610 2.543/2.413 31.268 1.762 0.847 19.866 0.626 2.910/2.937 41.200 4.445 0.839
EDTalk [2] 21.628 0.722 1.537/1.290 17.698 8.115 0.715 25.156 0.811 1.676/1.315 13.785 7.642 0.722
@@
TABLE I: A quantitative comparison with the SOTA methods on MEAD and HDTF. The highest scores for each metric are highlighted in bold. The symbols ’↑’ and ’↓’ denote that higher and lower values, respectively, indicate better performance.
Following previous work [40, 41], we adopt established metrics such as PSNR [60], SSIM [61], FID [62], E-FID [41], and FVD [63] to evaluate the visual quality and generative performance of our model. To assess facial motion accuracy and temporal consistency, we include Landmark Distance (LMD) [64] and the proposed Audio-Driven Facial Dynamics (ADFD) score. The SyncNet error metric is used to measure audio-visual synchronization, which is essential for evaluating lip-sync accuracy in talking face generation. In addition, we use CLIP-T [65], DINO [66], and Face Similarity [48] to assess perceptual relevance and identity preservation. CLIP-T evaluates consistency with the input prompt, DINO captures visual structure, and Face Similarity measures how closely the generated face matches the reference identity.
IV-B Implementation Details
Our framework, built upon Stable Diffusion v1.5 [67], leverages the $RealisticVision$ weights to fine-tune the model for audio-to-talking face generation. The implementation is carried out in PyTorch. We train our model using the AdamW [68] optimizer with a batch size of $32$ and a learning rate of 5e65e-6. The training process encompasses a multi-stage approach. Initially, each component is trained independently. IdentityNet is finetuned for 10 epochs, followed by the training of AnimateNet (excluding temporal attention) for $100$ epochs. Subsequently, in the second stage, the entire framework undergoes integrated training for $50$ epochs. All the training and testing phases are conducted on a single NVIDIA A100 GPU. The overall training takes approximately two weeks.
IV-C Quantitative Results
@@
To assess our model, we conducted a quantitative comparison with the state-of-the-art methods in audio-to-talking-face generation. The results are reported in Table I. PortraitTalk consistently outperforms the existing approaches across nearly all the metrics on the HDTF and MEAD datasets. Wav2Lip [12] achieves a relatively higher SyncNet score on HDTF, primarily because it utilizes this metric as one of its training loss functions. Our model achieves a satisfactory SyncNet score, surpassing Wav2Lip on the MEAD dataset, and excels in audio-lip alignment and facial structure accuracy, as indicated by LMD and ADFD metrics. Although EAMM has the highest ADFD score, our model ranks closely behind on the HDTF dataset, showing similar effectiveness. The slightly larger gap on the MEAD dataset is likely due to its wider range of emotional expressions and varying intensities, which directly impact lip movements.
@@
Figure 4: Qualitative comparison with the existing talking face generation methods. The results demonstrate that PortraitTalk surpasses previous methods in audio-lip alignment, identity resemblance, and expressiveness. Please note that the methods in the dash box use external emotion labels or reference videos to generate expressive videos.
Figure 5: Qualitative comparison of ablated variants of PortraitTalk, illustrating the effect of different model components. Each column shows output frames generated by models with specific components removed or added. These visualizations demonstrate the contribution of each design choice to the final visual quality and identity consistency of the generated talking head videos. This ablation study highlights how each component contributes to enhancing realism, expressiveness, and alignment with user intent.
IV-D Qualitative Results
@@
In Fig. 4, we compare the generation quality of PortraitTalk against the state-of-the-art approaches. MakeItTalk and Wav2Lip produce low-quality outputs with artifacts, notably around the mouth, which fail to blend seamlessly with adjacent facial features. Audio2Head [33], TalkLip [59], and IP-LAP [35] exhibit shortcomings in achieving precise audio-lip synchronization, with generated mouth movements often slightly open and displaying distortions across frames, thereby diminishing the realism of the output.
@@
In the emotional talking face generation, existing methods typically rely on discrete emotional labels or facial expressions from user-provided reference videos, which come with notable limitations. For instance, EAMM [13] and EmoGen [42] struggle with identity consistency, as shown in Fig. 4, where EAMM fails for the male subject and EmoGen for the female subject. Additionally, EmoGen restricts modifications to a facial square region, resulting in blurred areas that compromise realism. While EDTalk [2] improves emotional generation, it still suffers from jitter and irregular mouth shapes in frames. In contrast, our model demonstrates superior performance in both emotion-agnostic and emotional talking face generation. It excels at preserving identity consistency, ensuring precise audio-lip synchronization, and maintaining consistent head movements throughout the video.
@@
TABLE II: Quantitative results of the ablation study evaluating the impact of different components in PortraitTalk. The results show that excluding key elements such as temporal attention, head placement guidance, and the face encoder leads to notable drops in performance, particularly in synchronization and identity preservation metrics. The full model consistently achieves the best performance across all metrics, confirming the effectiveness of our complete system design.
IV-E Ablation Study
@@
Unified Attention Block: Initially, we merge the separate cross-attention blocks of IdentityNet into a single attention block to assess its impact on performance. The results are shown in Fig. 5 and Table II. The findings indicate a significant degradation in the model’s ability to preserve the intended identity, with its recognition capability reduced to identifying only the gender from text prompts.
@@
Impact of the Face Encoder: Further, we substitute our dedicated face encoder with the CLIP image encoder to explore the influence of the encoding mechanism on the generation capability of PortraitTalk. While there is an observable enhancement in image fidelity and identity consistency compared to the initial ablation scenario, the model still struggles to maintain detailed identity attributes across frames accurately. These experiments validate our decision to employ a specific face encoder optimized for facial feature extraction, underscoring its importance in achieving high-quality and consistent talking face generation.
Metric/Method FID↓ FVD↓ SSIM↑ E-FID↓
Aniportrait [38] 53.72 1042.93 0.731 1.94
V-Express [39] 59.34 1173.25 0.716 1.81
Hallo [40] 37.81 583.07 0.768 1.73
Echomimic [41] 32.51 536.41 0.775 1.68
PortraitTalk 31.86 586.19 0.786 1.64
TABLE III: Performance comparison of audio-driven talking face generation models using standard visual and temporal metrics.
@@
Number of Reference Images: We investigate the impact of using multiple reference images on the quality and realism of generated visual content. As demonstrated in Fig. 5, employing multiple images of the same identity, showcasing various expressions or head angles, significantly enhances the model’s capability to learn the underlying facial structure of a person. This strategy is essential for facilitating customization in generated talking face videos, enabling the model to respond to a variety of customization prompts effectively. In contrast, reliance on a single reference image leads to a considerable reduction in both the quality and realism of the generated faces. This issue is especially evident in the male case (Fig. 5), where the model produces an overly oval head shape.
@@
Impact of Structure and Temporal Attention: We also try to remove the head placement guidance and temporal attention mechanisms from AnimateNet. This modification leads to a noticeable decrease in coherence across the generated video frames. As illustrated in Fig. 5, while the lip synchronization remains somewhat accurate, there is a clear inconsistency in head placement and noticeable distortions between consecutive frames. Interestingly, without temporal attention, the generative model often fails to maintain consistency in the clothing of the generated talking faces. This variation underscores the crucial role of head placement guidance and temporal attention in maintaining the stability and continuity of head movements throughout the video.
Figure 6: Qualitative impact of mask loss on talking face generation. When mask loss is not applied during training, the model is more prone to generating artifacts such as text-like distortions and structural inconsistencies. Incorporating mask loss encourages the network to learn visual consistency by recovering corrupted regions, resulting in cleaner, more stable, and visually coherent frames.
IV-F Comparison with Recent Diffusion-based models
We compare PortraitTalk against recent diffusion-based audio-driven talking face generation frameworks, including AniPortrait [38], V-Express [39], Hallo [40], and EchoMimic [41]. As shown in Table III, while EchoMimic achieves the lowest FVD score, indicating stronger temporal consistency, PortraitTalk demonstrates a more balanced performance across all evaluation metrics. In particular, PortraitTalk obtained superior results in visual fidelity (FID, SSIM) and expression fidelity (E-FID), while maintaining competitive temporal coherence, leading to consistent identity preservation and realistic motion. It is also worth noting that PortraitTalk provides the flexibility to customize reference identity attributes, style, and the overall scene during generation, a capability not supported by current state-of-the-art methods. Visual examples of this functionality are presented in Sections IV-H and IV-I.
Method CLIP-T% ↑ DINO% ↑ Face.sim% ↑ SyncNet score↑
w/o mask loss 20.53 57.8 56.2 6.51
w mask loss 20.54 64.1 63.5 8.36
TABLE IV: Evaluation of the impact of mask loss on talking face generation. Results are reported across perceptual alignment (CLIP-T), structural consistency (DINO), identity preservation (Face Similarity), and audio-visual synchronization (SyncNet).
Figure 7: Balancing the contributions of reference images and text prompts in PortraitTalk. Our model demonstrates robustness in handling conflicting inputs, such as when the visual identity in the reference image does not align with the subject described in the prompt. In these cases, PortraitTalk adapts the generated face to reflect the textual description while preserving structural coherence from reference images. Additionally, by using multiple references, including from different identities, the model can blend features to synthesize coherent and novel outputs.
IV-G Impact of The Mask Loss
In this section, we evaluate the impact of incorporating mask loss during the training of PortraitTalk. As shown in Table IV, applying mask loss improves the overall generation quality, particularly in terms of identity preservation, visual fidelity, and audio-lip synchronization. By learning to recover corrupted regions, the model implicitly strengthens its understanding of visual consistency, leading to reduced artifacts and more stable frame-wise outputs. As illustrated in Fig. 6, this approach also mitigates unintended artifacts such as text-like distortions or detail inconsistencies issues often caused by the inherited biases of the underlying diffusion model.
IV-H Prompt and Reference Conflicts
PortraitTalk provides flexible control over both visual identity and text-driven attributes during generation. As illustrated in Fig. 7, PortraitTalk remains robust even when the reference image and prompt convey conflicting information. For example, when provided with a female reference image and a prompt describing a male subject, the model adapts facial features accordingly, generating outputs that reflect the textual description while maintaining structural consistency from the reference images. Furthermore, the multi-reference design offers additional flexibility. By supplying reference images from multiple identities, including across genders, the model can blend facial features to synthesize new, coherent identities. This demonstrates PortraitTalk’s ability to generalize beyond one-to-one mappings and to generate diverse, controllable outputs based on cross-modal and multi-source inputs.
Figure 8: Examples of diverse, expressive, and customizable talking faces generated by PortraitTalk. The model supports a wide range of applications, including identity interpolation, facial expression editing, and the modification of visual attributes such as age, gender, hairstyle, and background. These results demonstrate the effectiveness of PortraitTalk’s multimodal conditioning in enabling fine-grained, user-controllable synthesis while maintaining visual coherence and emotional expressiveness.
IV-I Applications and Limitations